OpenStudy (anonymous):

How do I "Find the vertex of the quadratic equation: y = 2x2 + 8x − 4."

OpenStudy (amistre64):

there are a few ways from very brunt mathing, to elequant techniques in form and design

OpenStudy (amistre64):

what methods are you aware of?

OpenStudy (anonymous):

None of the above? how about we go with the easiest way to do it and remember it?

OpenStudy (amistre64):

easiest way is to remember the general form of a quadratic as: \[ax^2+bx+c\] and remember that the value of the x component of the vertex is:\[x=-\frac{b}{2a}\]

OpenStudy (amistre64):

the y component is just inserting the value of the x component into the equation and solving

OpenStudy (amistre64):

y = 2x2 + 8x − 4 x = -8/(2*2) = -2 y = 2(-2)2 + 8(-2) − 4 vertex: (-2, f(-2))

OpenStudy (anonymous):

f?

OpenStudy (anonymous):

\(y=f(x)=2x^2+8x-4\)

OpenStudy (amistre64):

y = f(x) y = f(-2) simply means the value of the function at x=-2

OpenStudy (anonymous):

I am so confused 0.0

OpenStudy (mertsj):

Find the x coordinate of the vertex by using x = -b/2a = -8/4=-2 Plug that back in the find the y coordinate: \[2(-2)^2+8(-2)-4=2(4)-16+4=8-16-4=-12\]